Payment of 400.00 every 5 months for 3 years that start at the end of 5 years. What is the period of deferral?


1
Expert's answer
2022-01-18T06:09:44-0500

period of deferral:

"k=nt-1"

where n is number of times compounded,

t is time


we have:

"nt=5\cdot12/5=12"


then:

"k=12-1=11" periods "=11\cdot 5=55" month